40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12085     Entry time: Thu Apr 21 14:25:52 2016     In reply to: 12083
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: Green light recovered 

I've made progress on the new layout up to the doubling oven. After doing the coarse alignment with the diode current to the NPRO at ~1A, I turned it back up to the nominal 2A. I then rotated the HWP before the IR Faraday such that only ~470mW of IR power is going into the doubler (the rest is being dumped on razor beam dumps). After tuning the alignment of the IR into the doubling oven using the steering mirror + 4 axis translation stage on which the doubling oven is mounted, I get ~3.2mW of green after the harmonic separator and a HR mirror for green. The mode looks pretty good to the eye (see attachment #1), and the conversion efficiency is ~1.45%/W - which is somewhat less than the expected 2%/W but in the ballpark. It may be that some fine tweaking of the alignment + polarization while monitoring the green power can improve the situation a little bit (I think it may go up to ~4mW, which would be pretty close to 2%/W conversion efficiency). The harmonic separator also seems to be reflecting quite a bit of green light along with IR (see attachment #2) - so I'm not sure how much of a correction that introduces to the conversion efficiency. 

While doing the alignment, I noticed that some amount of IR light is actually transmitted through the HR mirrors. With ~500mW of incident light at ~45 degrees, this transmitted light amounts to ~2mW. Turns out that this is also polarization dependant (see attachment #3) - for S polarized light, as at the first two steering mirrors after the NPRO, there is no transmitted light, while for P-polarized light, which is what we want for the doubling crystal, the amount transmitted is ~0.5%. The point is, I think the measured levels are consistent with the CVI datasheet. We just have to take care find all these stray beams and dump them.

I will try and optimize the amount of green power we can get out of the doubler a little more (but anyway 3mW should still be plenty for ALS). Once I'm happy with that, I will proceed with laying out the optics for mode-matching the green to the arm.

Attachment 1: IMG_6567.JPG  1.031 MB  | Show | Show all
Attachment 2: IMG_6568.JPG  1.438 MB  | Show | Show all
Attachment 3: CVI_reflectivity.jpeg  71 kB  | Show | Show all
ELOG V3.1.3-